A car has a gas mileage of 12 km/L. Convert this rating to miles per gallon


1
Expert's answer
2021-02-01T03:59:35-0500

"\frac{12km}{1L}\times\frac{1mi}{1.609km}\times\frac{3.785L}{1gal}=28mi/gal"


Answer: 28 mi/gal